OpenStudy (anonymous):

can someone help me please?! how does one find tan 13pi/4 without using a calculator or decimals?

OpenStudy (campbell_st):

remember the fact that \[revolution = 2\pi\] \[\frac {13\pi}{4} = 3 \frac{1}{4} \pi\] so the angle is pi/4 more that 1 and 1/2 revolutions so its a 3rd a quadrant angle where tan is positive so \[\tan(\frac{13\pi}{4} = \tan (\frac{\pi}{4})\] should be easy to find now

OpenStudy (anonymous):

Thank you so much! Would 1 be the correct answer? @campbell_st

OpenStudy (campbell_st):

that would be my answer